### v2

# Quasi-Clique Discovery via Energy Diffusion

Abstract
Discovering quasi-cliques—subgraphs whose edge density exceeds a given threshold—is a fundamental task in graph mining with applications to web spam detection, fraud screening, and e-commerce recommendation. However, existing methods for quasi-clique discovery on large-scale web graphs are often sensitive to random seeds or lack of explicit edge-density guarantees, making the task challenging in practice. This paper presents EDQC, an energy diffusion–based method for quasi-clique discovery. EDQC first employs an adaptive energy diffusion process to generate an energy ranking that highlights structurally cohesive regions. Guided by this energy ranking, the algorithm identifies a high-quality subgraph by minimizing conductance, a standard measure from community detection. This subgraph is then refined to meet the specified density threshold. Extensive experiments on 75 real-world graphs show that EDQC finds larger quasi-cliques on most datasets, with consistently lower variance across runs and competitive runtime. To the best of our knowledge, EDQC is the first method to incorporate energy diffusion into quasi-clique discovery.
1 Introduction
Dense subgraph models are fundamental to graph mining on large-scale web graphs, enabling web spam detection gibson2005discovering ; lin2021mining , fraud screening zheng2019author ; fang2022densest , and e-commerce recommendation schafer2001commerce ; song2021subgraph . Among these models, the clique represents the most stringent form, requiring that every pair of vertices be directly connected to each other saito2007large . However, due to noise and missing links in real-world graphs, exact cliques are rarely observed. To relax this strict requirement, several approximate dense subgraph models have been proposed, including quasi-cliques abello2002massive ; pei2005mining , $k$-clubs mokken1979cliques , and $k$-plexes seidman1978graph . In particular, the $\gamma$-quasi-clique, defined by a minimum edge-density threshold, is one of the most widely adopted formulations chen2021nuqclq ; pang2024similarity ; liu2024optimization .
Given an undirected graph $G$ and an edge-density threshold $\gamma\in(0,1]$, the Maximum Quasi-Clique Problem (MQCP) seeks the largest vertex set $S$ such that the induced subgraph $G[S]$ exhibits an edge density of at least $\gamma$. This edge-density criterion provides a more direct measure of structural cohesion compared to the degree-based definitions utilized by methods such as FastQC konar2024optimal and IterQC liu2008effective . However, the MQCP is NP-hard even for fixed values of $\gamma$ pattillo2013maximum and does not possess the hereditary property that allows for efficient pruning in problems related to cliques xia2025maximum . Nonetheless, quasi-clique discovery remains highly valuable, as real-world graphs typically exhibit dense yet incomplete connectivity rather than perfect cliques, making quasi-cliques crucial for identifying significant or anomalous substructures.
Figure 1: A clique and a 0.8-quasi-clique.
One practical application of quasi-clique discovery is identifying link farm spam pages in web graphs pang2024similarity , where densely connected spam pages form $\gamma$-quasi-cliques, enabling the identification of suspicious link farms even under noise and missing links. As shown in Figure 1, the dashed edges in the left subgraph form a size-4 clique, representing a fully connected link farm. In the right subgraph, the dashed edges form a 0.8-quasi-clique, still suspicious despite one missing link, capturing a potential link farm under partial observation or evasion.
Despite recent progress, discovering quasi-cliques in real large-scale web graphs remains challenging due to both scale and combinatorial complexity. While exact algorithms for MQCP, typically based on integer programming or techniques such as branch-and-bound uno2010efficient ; pastukhov2018maximum ; ribeiro2019exact ; rahman2024fast , guarantee optimal solutions, their exponential time complexity restricts their applicability to large graphs. As a practical alternative, numerous heuristic methods have been proposed to enable scalable quasi-clique discovery, which can be roughly grouped into four categories. First, greedy expansion techniques, such as the greedy strategy with local refinement introduced in tsourakakis2013denser and the restart-enabled search in RIG oliveira2013construction , prioritize speed but are vulnerable to becoming trapped in local optima. Second, metaheuristic frameworks like TSQC djeddi2019extension and OBMA zhou2020opposition employ tabu search and opposition-based learning, offering enhanced robustness at a higher computational cost. Third, local search algorithms, including NuQClq chen2021nuqclq and its variants NuQClqF1 and NuQClqR1 liu2024optimization , utilize bounded configuration checking and adaptive rules for candidate refinement. However, despite their sophisticated local optimizations, these methods often exhibit instability due to their dependence on random seeds. Lastly, similarity-based heuristics such as NBSim and its accelerated variant, FastNBSim, utilize vertex neighborhood overlap to identify and expand quasi-cliques pang2024similarity . While both methods are highly efficient, FastNBSim introduces MinHash-based approximations, making it sensitive to random seeds and potentially less stable. Moreover, neither method supports a user-specified edge-density threshold. Instead, the edge-density of the returned subgraph emerges implicitly from similarity parameters. Consequently, there is no guarantee that the output meets a specified minimum edge-density requirement. This limits their suitability for applications that require strict control over quasi-clique edge-density thresholds.
Among these four categories, local search and similarity-based methods have been the most competitive for discovering large quasi-cliques. The state-of-the-art algorithms for MQCP include NBSim and its accelerated variant, FastNBSim pang2024similarity , as well as the local search algorithms NuQClqF1 and NuQClqR1 liu2024optimization . However, these methods remain sensitive to random seeds or lack of explicit edge-density guarantees.
Energy diffusion is a natural candidate for addressing the challenges above. Diffusion methods such as PageRank page1999pagerank and the heat kernel chung2007heat have shown strong performance on structure-aware tasks, e.g., vertex ranking and community detection kloster2014heat ; gleich2015multilinear , where diffused mass tends to localize within cohesive regions. These tasks overlap with quasi-clique discovery in highlighting tightly interlinked vertices, yet they are not equivalent because quasi-cliques must satisfy a specified minimum edge-density on induced subgraphs. This connection motivates exploring energy diffusion for quasi-clique discovery to reduce sensitivity to random seeds in practice and to furnish a stable candidate region for density-constrained extraction.
This paper introduces EDQC, an energy diffusion-based method for quasi-clique discovery. EDQC begins with an adaptive energy diffusion phase, where the energy transfer rate is dynamically adjusted based on the local energy distribution. This phase runs for a fixed number of iterations from each source vertex, generating an energy ranking that highlights structurally cohesive regions. Guided by this energy ranking, EDQC then employs a principled extraction mechanism from community detection theory. It applies a sweep cut to the energy-ranked vertices to identify a high-quality subgraph that minimizes conductance. This subgraph is subsequently refined through a structure-driven procedure to guarantee it meets the specified density threshold. Extensive experiments on 75 real-world graphs, including classic web graphs, demonstrate that EDQC discovers larger quasi-cliques on most datasets, with consistently lower variance and competitive runtime. To the best of our knowledge, EDQC is the first method to incorporate energy diffusion into quasi-clique discovery.
2 Preliminaries
All graphs considered in this paper are undirected and simple (no loops or parallel edges). Let $G=(V,E)$ represent a graph with a vertex set $V$ and an edge set $E$. Two vertices $u$ and $v$ within $V$ are termed adjacent if $uv\in E$, and one vertex is referred to as a neighbor of the other. The set of all neighbors for a vertex $u$ is known as its neighborhood, denoted $N(u)$, while the closed neighborhood of $u$ is denoted $N[u]=N(u)\cup\{u\}$. The degree of a vertex $u$, denoted as $d(u)$, indicates the number of its neighbors, i.e., $d(u)=|N(u)|$.
Let $n=|V|$ and $m=|E|$. For a vertex set $S\subseteq V$, the induced subgraph is $G[S]$ with vertex set $S$ and edge set $E(S)=\{\{u,v\}\in E\mid u,v\in S\}$. Let $m(S)=|E(S)|$ denote the number of internal edges in $G[S]$, and define the internal degree of $u\in S$ by
$d_{S}(u)=|N(u)\cap S|.$
The edge density of $S$ (and of $G[S]$) is
$\eta(S)=\begin{cases}\dfrac{2\,m(S)}{|S|(|S|-1)},&|S|\geq 2,\\[6.0pt] 0,&otherwise,\end{cases}$
and $\eta(G)$ is shorthand for $\eta(V)$. Throughout this paper, unless otherwise stated, density means edge density.
Definition 1 ($\gamma$-Quasi-Clique).
Given $G=(V,E)$ and a density threshold $\gamma\in(0,1]$. A vertex set $S\subseteq V$ induces a $\gamma$-quasi-clique if the induced subgraph $G[S]$ has edge density at least $\gamma$, i.e.,
$\eta(G[S])\;\geq\;\gamma.$
For brevity, we also refer to $G[S]$ itself as a $\gamma$-quasi-clique. A $\gamma$-quasi-clique $G[S]$ is maximal if there is no $v\in V\setminus S$ such that $\eta\!\left(G[S\cup\{v\}]\right)\geq\gamma$.
Definition 2 (Maximum Quasi-Clique Problem).
Given $G=(V,E)$ and a density threshold $\gamma\in(0,1]$, the Maximum Quasi-Clique Problem (MQCP) seeks
$S^{*}\in\arg\max_{S\subseteq V}\ \Big\{|S|\ \Big|\ \eta(G[S])\geq\gamma\Big\}.$
The induced subgraph $G[S^{*}]$ is a maximum $\gamma$-quasi-clique. When $\gamma=1$, MQCP reduces to the classical Maximum Clique Problem. For any fixed $\gamma$, MQCP is NP-hard and may admit multiple optimal solutions pattillo2013maximum .
Definition 3 (Conductance).
Conductance is a standard measure used to evaluate how well a vertex set forms a community andersen2006local . For a non-empty, proper subset of vertices $S\subset V$, its volume is defined as $\mathrm{vol}(S)=\sum_{u\in S}d(u)$, and the cut between $S$ and its complement is $\mathrm{cut}(S,V\setminus S)=|\{\{u,v\}\in E\mid u\in S,v\in V\setminus S\}|$. The conductance of $S$ is then:
$\Phi(S)=\frac{\mathrm{cut}(S,V\setminus S)}{\min(\mathrm{vol}(S),\mathrm{vol}(V\setminus S))}.$
A lower conductance value indicates a more clearly defined community, signifying few edges connecting the set to the rest of the graph relative to its internal connectivity.
3 The EDQC Algorithm
This section details the proposed Energy Diffusion-based Quasi-Clique discovery (EDQC) algorithm. Given a graph $G=(V,E)$ and a density threshold $\gamma$, EDQC identifies a large vertex set $S$ that induces a $\gamma$-quasi-clique. The overall pipeline is outlined in Algorithm 1.
The algorithm iterates through vertices, treating each as a source to initiate a two-stage process. First, an adaptive energy diffusion process generates an energy distribution over the graph. Second, guided by this energy distribution, a principled extraction mechanism identifies a high-quality quasi-clique. EDQC maintains the largest quasi-clique found across all source vertices. Processing vertices in non-increasing degree order is an efficiency heuristic rossi2014fast , as it ensures high-degree vertices are processed as energy sources first, thereby increasing the likelihood of finding a large quasi-clique early in the search.
Input: Graph $G=(V,E)$, density threshold $\gamma$, diffusion rounds $T$, activation threshold $\theta$
Output: Vertex set $S^{*}$ of a large $\gamma$-quasi-clique
1 $S^{*}\leftarrow\emptyset$;
2 foreach $v\in V$ in non-increasing order of degree do
[…]
5 if $|S|>|S^{*}|$ then
6    $S^{*}\leftarrow S$;
7
8
return $S^{*}$
[…]
3.1 Adaptive Energy Diffusion
The first stage, detailed in Algorithm 2, simulates the diffusion of energy from a source vertex $v$ over $T$ synchronous rounds. The goal is to produce an energy function $f:V\to\mathbb{R}_{\geq 0}$ where energy is concentrated in structurally cohesive regions surrounding the source.
The diffusion process is designed to be adaptive. The fraction of energy a vertex releases depends on its current energy relative to its neighborhood, a dynamic governed by an adaptive fraction $\alpha(u)$. This design is deliberate and distinct from standard diffusion mechanisms such as the heat kernel chung2007heat or PageRank page1999pagerank , which often treat connections uniformly. Our adaptive fraction makes the diffusion process sensitive to the evolving energy landscape, enabling EDQC to be more conservative in sparse regions and more aggressive within dense ones.
The diffusion starts by initializing a unit energy at the source vertex $v$ (line 1). In each of the $T$ synchronous rounds (lines 2–14), the algorithm first identifies the set of active vertices, $A$, based on the activation threshold $\theta$ (line 3). To ensure updates are synchronous, a temporary copy of the energy function, $f^{\prime}$, is created (line 4). The diffusion rule is then applied to each active vertex $u\in A$ with a non-zero degree ($d(u)>0$). Isolated vertices are skipped as they have no neighbors to which energy can be transferred, a condition necessary for the conservation of total energy. For each such vertex $u\in A$ (line 5), the algorithm computes its adaptive fraction $\alpha(u)$ and the total energy to transfer, $\delta(u)$ (lines 7–8). This energy is then stochastically allocated among its neighbors using weights $\phi_{uw}$ derived from an exponential distribution (lines 9–10). Subsequently, the energies of the active vertex and its neighbors are updated in the temporary function $f^{\prime}$ (lines 11–13). After all active vertices have been processed, the energy function $f$ is updated with the temporary copy, completing the round (line 14).
Input: Graph $G=(V,E)$, source vertex $v$, diffusion rounds $T$, activation threshold $\theta$
Output: Energy function $f:V\to\mathbb{R}_{\geq 0}$
1 Initialize $f(v)\leftarrow 1$ and $f(x)\leftarrow 0$ for $x\neq v$;
2 for $t=1$ to $T$ do
3 $A\leftarrow\{x\in V\mid f(x)>\theta\}$;
4 $f^{\prime}\leftarrow f$;
5 foreach $u\in A$ do
6    if $d(u)>0$ then
7       $\alpha(u)\leftarrow\frac{\sum_{w\in N(u)}f(w)+f(u)}{\sum_{w\in N(u)}f(w)+2f(u)}$;
8       $\delta(u)\leftarrow\alpha(u)f(u)$;
9         Sample $\xi_{uw}\sim\mathrm{Exp}(1)$ for each $w\in N(u)$;
10       $\phi_{uw}\leftarrow\frac{\xi_{uw}}{\sum_{x\in N(u)}\xi_{ux}}$ for each $w\in N(u)$;
11       $f^{\prime}(u)\leftarrow f^{\prime}(u)-\delta(u)$;
12       foreach $w\in N(u)$ do
13          $f^{\prime}(w)\leftarrow f^{\prime}(w)+\delta(u)\phi_{uw}$;
14
15
16
17 $f\leftarrow f^{\prime}$;
18
return $f$
[…]
Properties.
This diffusion process has several key properties.
Proposition 1 (Energy Conservation).
The total energy is conserved in each round: $\sum_{x\in V}f^{\prime}(x)=\sum_{x\in V}f(x)$.
Proof.
Consider a fixed round and an arbitrary active vertex $u$ where $d(u)>0$. The synchronous updates associated with $u$ results in a change of $\delta(u)$ for $u$ and an addition of $\delta(u)\phi_{uw}$ for each neighbor $w\in N(u)$, satisfying $\sum_{w\in N(u)}\phi_{uw}=1$. Therefore, $\delta(u)=\sum_{w\in N(u)}\delta(u)\phi_{uw}$. Summing over all active vertices leads to the conclusion that $\sum_{x\in V}f^{\prime}(x)=\sum_{x\in V}f(x)$. Here, we should note that the energy of vertices for which $d(u)=0$ or $f(u)=0$ remains unchanged in the energy diffusion process. ∎
The conservation of energy is a crucial property for the algorithm’s stability. It guarantees that the process only redistributes the initial energy, rather than arbitrarily creating or losing it, thus ensuring a fair basis for comparing the energy distributions originating from different source vertices.
Proposition 2 (Properties of $\alpha(u)$).
For an active vertex $u$ with $f(u)>0$ and $d(u)>0$, the adaptive fraction $\alpha(u)\in[1/2,1)$ and is monotonically increasing with the total energy of its neighbors, $S(u)=\sum_{w\in N(u)}f(w)$.
Proof.
We begin by rewriting the formula for $\alpha(u)$ (line 7, Algorithm 2) as $\alpha(u)=1-\frac{f(u)}{S(u)+2f(u)}$. Since $S(u)\geq 0$ and $f(u)>0$ for an active vertex, the denominator $S(u)+2f(u)$ is strictly positive. The fractional term is bounded by $0<\frac{f(u)}{S(u)+2f(u)}\leq\frac{f(u)}{2f(u)}=1/2$. Consequently, $\alpha(u)$ is bounded by $1-1/2\leq\alpha(u)<1-0$, which gives $\alpha(u)\in[1/2,1)$. For monotonicity, the partial derivative with respect to $S(u)$ is:
$\frac{\partial\alpha(u)}{\partial S(u)}=\frac{f(u)}{(S(u)+2f(u))^{2}}>0.$
This confirms that $\alpha(u)$ is strictly increasing with $S(u)$. ∎
The adaptive nature of our diffusion is formally established by this proposition. The bounds on $\alpha(u)$ guarantee a process that is both consistently active and stable, preventing the diffusion from stalling. Concurrently, the monotonicity is the key mechanism driving the algorithm’s core premise: it mathematically ensures that energy is exchanged more readily within already high-energy (and thus likely dense) regions, accelerating the concentration of energy.
Proposition 3 (Unbiased Distribution in Expectation).
The expected fraction of energy that an active vertex $u$ transfers to a specific neighbor $w\in N(u)$ is uniform across all neighbors: $\mathbb{E}[\phi_{uw}]=1/d(u)$.
Proof.
The weights $\phi_{uw}$ are derived by normalizing i.i.d. exponential random variables $(\xi_{uw})_{w\in N(u)}$. This construction is a standard method for generating a random point from a uniform distribution on the standard simplex. By symmetry, the expectation of each component is identical. Since $\sum_{w\in N(u)}\phi_{uw}=1$, by linearity of expectation, $\sum_{w\in N(u)}\mathbb{E}[\phi_{uw}]=1$. As there are $d(u)$ identical terms in the sum, it follows that $d(u)\cdot\mathbb{E}[\phi_{uw}]=1$, which proves $\mathbb{E}[\phi_{uw}]=1/d(u)$. ∎
While the exponential sampling introduces stochasticity to explore different structural paths, this proposition guarantees that the process remains unbiased in expectation. On average, all neighbors are treated fairly, ensuring that the algorithm itself introduces no arbitrary structural bias during the diffusion.
Theorem 1 (Complexity of Algorithm 2).
The EnergyDiffusion algorithm runs in $O(T(n+m))$ time.
Proof.
The algorithm performs $T$ rounds. We analyze the computational cost of a single round. First, identifying the active set $A$ requires a full scan of all vertices, which takes $O(n)$ time. Next, the algorithm iterates through each active vertex $u\in A$. For each such vertex, all calculations—including the summations for $\alpha(u)$ and $\phi_{uw}$, and the subsequent updates—are proportional to its degree, $d(u)$. Therefore, the total work for processing all active vertices in a round is bounded by the sum of degrees of all vertices in the graph, which is $\sum_{u\in V}O(d(u))=O(m)$. Summing the costs within a single round gives a complexity of $O(n+m)$. As the process repeats for $T$ rounds, the total complexity of the algorithm is $O(T(n+m))$. ∎
3.2 Quasi-Clique Extraction
The second stage of EDQC takes the final energy function $f$ returned by Algorithm 2 and produces a vertex set $S$ that induces a $\gamma$-quasi-clique. This extraction process (Algorithm 3) has three main operations: an initial ranking of vertices, the identification of a high-quality initial subgraph, and a refinement procedure.
First, the algorithm identifies the set of active vertices $A$ and sorts them in descending order of their energy scores to create an ordered sequence $P$ (lines 1–2). This ranking is crucial, as it places vertices most likely to belong to a cohesive community at the forefront.
Next, to identify a high-quality initial candidate set, the algorithm performs a sweep cut andersen2006local on the energy-ranked list $P$ (line 3). This procedure examines the nested prefixes of the list and selects the one that minimizes conductance, as it represents the most well-defined community structure. This strategy is highly effective because the initial energy ranking provides a high-quality vertex permutation. It is a known principle in local graph clustering that sweep cuts perform best when applied to an ordering that places vertices from the same community contiguously andersen2006local . Our energy diffusion process is designed for precisely this purpose: it creates an ordering where structurally cohesive vertices are naturally placed at the top. This, in turn, enables the sweep cut to accurately identify the boundary of the best candidate subgraph.
Finally, this initial set $S$ undergoes a refinement procedure (lines 4–15). The refinement begins with a pruning loop (lines 4–6). If the initial set $S$ does not meet the density threshold, this loop iteratively removes the vertex with the minimum internal degree $d_{S}(u)$ until the set $S$ induces a valid $\gamma$-quasi-clique. Following this, an expansion phase seeks to enlarge the set (lines 8–15). In each pass, the remaining candidates $R$ are re-sorted based on their connectivity to the current set $S$ (line 10). The algorithm then iterates through this sorted list $R$, attempting to greedily add the first vertex $v$ that maintains the density constraint (lines 12–13). Upon a successful insertion, the current iteration over the candidate list is immediately terminated (line 15), allowing the algorithm to restart the process with an up-to-date candidate ranking in the next iteration.
Input: Graph $G=(V,E)$, energy function $f$, density threshold $\gamma$, activation threshold $\theta$
Output: A vertex set $S$ that induces a $\gamma$-quasi-clique
1
[…]
3 $P\leftarrow$ vertices of AA sorted by decreasing $f$;
4
[…]
6
7while $|S|>1$ and $\eta(G[S])<\gamma$ do
8 $u\leftarrow\arg\min_{x\in S}d_{S}(x)$;
9 $S\leftarrow S\setminus\{u\}$;
10
11
[…]
13 while changed do
14 $\textit{changed}\leftarrow\textbf{false}$;
15 $R\leftarrow$ vertices of $P\setminus S$, sorted by connections to $S$;
16 foreach $v\in R$ do
17    if $\eta(G[S\cup\{v\}])\geq\gamma$ then
18       $S\leftarrow S\cup\{v\}$;
19       $\textit{changed}\leftarrow\textbf{true}$;
20       break;
21
22
23
return $S$
[…]
Theorem 2 (Complexity of Algorithm 3).
Let $k=|A|$ be the number of active vertices. The ExtractQuasiClique stage runs in $O(k^{2}+m(A))$, where $m(A)$ is the number of edges in the induced subgraph $G[A]$.
Proof.
The overall complexity is determined by the main steps of the algorithm. Sorting the $k$ active vertices takes $O(k\log k)$ time. The subsequent sweep cut, which involves calculating conductance for $k$ prefixes, can be implemented in $O(m(A))$ time with incremental updates for the cut and volume. The final refinement stage is the most computationally intensive. The pruning and expansion loops can, in the worst case, lead to a series of $O(k)$ vertex additions or removals. With incremental maintenance of internal degrees and edge counts, each step of these loops can take up to $O(k)$ time to find the best candidate to remove or to scan through remaining candidates for addition. This results in a worst-case complexity of $O(k^{2})$ for the refinement phase. Summing these costs, the total complexity is $O(k\log k+m(A)+k^{2})$, which simplifies to $O(k^{2}+m(A))$. ∎
Theorem 3 (Complexity of EDQC).
The theoretical worst-case time complexity of the EDQC algorithm (Algorithm 1) is $O(n^{3})$.
Proof.
The algorithm’s main loop iterates up to $n$ times. Each iteration calls two procedures: EnergyDiffusion, with a complexity of $O(T(n+m))$, and ExtractQuasiClique, with a worst-case complexity of $O(n^{2}+m)$. As $T$ is a small constant, the cost of a single iteration is dominated by the extraction step, at $O(n^{2}+m)$. Therefore, the total worst-case complexity is $O(n^{3}+nm)$. Since $m=O(n^{2})$ for any simple graph, this simplifies to $O(n^{3})$. The initial sorting of vertices by non-increasing degree is a lower-order term that does not affect this bound.
We note that this is a theoretical worst-case analysis. In practice, EDQC is run as a time-bounded heuristic, and its empirical efficiency is demonstrated in our experimental section. ∎
4 Experimental Evaluation
We conducted a comprehensive experimental evaluation to validate the performance and robustness of EDQC. This section is organized as follows. We first detail the experimental setup, the baseline methods, and the datasets used. We then present the main results on effectiveness and efficiency, followed by a statistical analysis to confirm the significance of our findings. Subsequently, we provide an in-depth analysis of the algorithm, including an ablation study to verify the contribution of its key components and an empirical study to validate its core premise. The section concludes with a summary of our parameter sensitivity analysis.
4.1 Experimental Setup
All methods were implemented in C++ and compiled with g++ 9.4.0 using the ‘-O3’ optimization flag. The experiments were conducted on a machine equipped with an AMD EPYC 7H12 CPU (1.5 GHz base frequency) and 128 GB of RAM, running Ubuntu 20.04.
Each baseline was run with the parameter settings recommended in its original paper. For EDQC, we fixed the diffusion rounds at $T=2$ and the activation threshold at $\theta=10^{-4}$. These values were selected based on a sensitivity study, the details of which are provided in Section 4.8.
Due to inherent randomness, all non-deterministic methods (all except NBSim) were executed 10 times with random seeds from 1 to 10; NBSim, being deterministic, was run once. For the non-deterministic methods, we report their average quasi-clique size and standard deviation (STD) over the 10 runs, presented in the format ‘average size $\pm$ STD’, where the standard deviation is reported only if it is non-zero. All runs were subject to a 60-second timeout. If an algorithm failed to find any valid $\gamma$-quasi-clique in all 10 runs for an instance, we mark its result as ‘N/A’. The 60-second timeout reflects common practice in recent time-bounded heuristic studies lin2017reduction ; zheng2023farsighted .
4.2 Baseline Methods
We compared EDQC against four state-of-the-art algorithms:
  • NBSim/FastNBSim pang2024similarity : Similarity-based heuristics that expand subgraphs based on neighborhood overlap. FastNBSim accelerates NBSim via MinHash-based approximation.
  • NuQClqF1/NuQClqR1 liu2024optimization : Advanced variants of NuQClq chen2021nuqclq . These algorithms employ information feedback models that integrate historical solutions using fitness-weighted scores to guide local search.
4.3 Datasets
We evaluated all algorithms on 75 real-world datasets from SNAP snapnets and the Network Repository nr . The collection covers diverse domains, including web, social, and communication graphs. The datasets span a wide spectrum of scales, with vertex counts ($n$) ranging from 62 to over 16.7 million, and edge densities spanning from $8.41\times 10^{-2}$ down to as low as $1.87\times 10^{-7}$. This diversity provided a challenging testbed for evaluating the scalability and effectiveness of the algorithms.
4.4 Main Results
We now present the primary results comparing the effectiveness and efficiency of EDQC against the baselines. A key challenge in this comparison is that NBSim and FastNBSim do not accept a density threshold $\gamma$ as input. To facilitate a fair evaluation, we designed two separate experimental settings (Table 1 and Table 2). To save space, we omit instances where all four compared algorithms found solutions of the same average size.
Table 1: Quasi-clique sizes with the threshold $\gamma$ set by NBSim’s output.
Instance $\gamma$ NBSim NuQClqF1 NuQClqR1 EDQC
rt-retweet 1 3 4 4 4
web-polblogs 1 5 9 9 9
email-Eu-core 1 12 18 18 18
ego-facebook 0.99 71 92 92 92
ca-Erdos992 1 6 8 8 8
socfb-CMU 0.98 20 53 53 53
ech-WHOIS 0.98 57 79 79 79
web-indochina-2004 1 50 49 $\pm$ 3.16 50 50
socfb-UCSB37 0.99 40 64 64 64
socfb-UConn 0.99 22 58 58 58
socfb-UCLA 0.79 19 96 96 96
socfb-Berkeley13 1 21 42 42 42
socfb-Wisconsin87 0.98 12 46 46 46
soc-epinions 1 10 16 16 16
Cit-HepTh 1 12 22.9 $\pm$ 0.32 23 23
socfb-Indiana 0.85 27 94 94 94
socfb-UIllinois 1 14 57 57 57
Cit-HepPh 0.98 14 23 23 23
socfb-UF 0.99 24 67 66.8 $\pm$ 0.63 67
Email-Enron 0.97 10 25 25 25
socfb-Penn94 0.98 14 56 55.3 $\pm$ 2.21 56
soc-brightkite 0.99 36 45 45 45
socfb-OR 0.6 19 108 108 108
soc-slashdot 1 12 26 26 26
witter_combined 0.98 79 89.3 $\pm$ 11.7 89.3 $\pm$ 11.7 93
G_n_pin_pout 1 3 4 4 4
yahoo-msg 0.99 22 20.6 $\pm$ 2.07 20.6 $\pm$ 2.01 23
soc-douban 1 5 11 11 11
wave 1 5 6 6 6
Loc-Gowalla 0.99 31 23.8 $\pm$ 2.53 23 31
soc-gowalla 0.99 31 23 23.8 $\pm$ 2.53 31
web-Stanford 0.99 67 64.2 $\pm$ 3.43 59.9 $\pm$ 4.12 67.5 $\pm$ 0.71
cnr-2000 0.99 89 84.8 $\pm$ 3.61 81.8 $\pm$ 0.63 89
web-NotreDame 1 155 154.3 $\pm$ 0.48 154.2 $\pm$ 0.42 155
ca-MathSciNet 1 25 24.6 $\pm$ 0.84 24.7 $\pm$ 0.67 25
coPapersDBLP 1 337 333.5 $\pm$ 4.12 333.5 $\pm$ 4.12 337
auto 1 6 7 7 7
web-it-2004 1 432 431.8 $\pm$ 0.42 432 432
soc-delicious 0.58 22 80 80 80
ca-coauthors-dblp 1 337 333.5 $\pm$ 4.12 333.5 $\pm$ 4.12 337
eu-2005 0.99 405 191.3 $\pm$ 115.84 127.2 $\pm$ 26.92 405.1 $\pm$ 0.32
web-Google 0.99 48 39.2 $\pm$ 7.86 39.5 $\pm$ 7.41 48
rgg_n_2_20_s0 0.57 17 41 41 40.2 $\pm$ 0.79
rgg_n_2_21_s0 1 18 18.1 $\pm$ 0.57 18.1 $\pm$ 0.57 19
rgg_n_2_22_s0 1 19 18.8 $\pm$ 0.79 18.6 $\pm$ 0.7 20
rgg_n_2_23_s0 0.99 19 18.3 $\pm$ 0.48 19 $\pm$ 1.41 22
rgg_n_2_24_s0 0.98 20 19.7 $\pm$ 0.95 18.7 $\pm$ 0.48 23
Table 2: Quasi-clique sizes with $\gamma$ set by FastNBSim’s max density over 10 runs. For FastNBSim, stats are on successful runs only (rate shown in parentheses).
Instance $\gamma$ FastNBSim NuQClqF1 NuQClqR1 EDQC
soc-dolphins 0.83 4 (4/10) 6 6 6
rt-retweet 1 3 (7/10) 4 4 4
ca-netscience 1 8.1 $\pm$ 0.88 (10/10) 9 9 9
web-polblogs 0.93 6 (2/10) 13 13 13
rt-twitter-copen 1 3.6 $\pm$ 0.53 (7/10) 4 4 4
email-Eu-core 0.94 21 (1/10) 29 29 29
web-edu 1 27.4 $\pm$ 3.1 (10/10) 30 30 30
ego-facebook 0.96 102 (1/10) 123 123 123
ca-GrQc 1 37.8 $\pm$ 3.2 (4/10) 44 44 44
web-spam 1 15.5 $\pm$ 0.71 (2/10) 20 20 20
ca-Erdos992 1 5 (6/10) 8 8 8
socfb-CMU 0.97 28 (1/10) 56 56 56
ech-WHOIS 0.96 73 (1/10) 91 91 91
ca-HepPh 1 202.7 $\pm$ 24.04 (10/10) 239 239 239
web-indochina-2004 1 49.9 $\pm$ 0.32 (10/10) 49 $\pm$ 3.16 50 50
web-BerkStan 1 26.7 $\pm$ 2.83 (10/10) 29 29 29
socfb-UCSB37 0.98 47 (1/10) 70 70 70
web-webbase-2001 1 32.9 $\pm$ 0.32 (10/10) 33 33 33
socfb-UConn 0.97 34 (1/10) 65 65 65
ca-AstroPh 1 51.8 $\pm$ 3.12 (10/10) 57 57 57
socfb-UCLA 0.98 49 (1/10) 63 63 63
ca-CondMat 1 22.63 $\pm$ 2.83 (8/10) 26 26 26
socfb-Berkeley13 0.95 29 (1/10) 62 62 62
socfb-Wisconsin87 0.91 34 (1/10) 65 65 65
soc-epinions 0.95 16 (1/10) 21 21 21
Cit-HepTh 0.92 26 (1/10) 41 41 41
socfb-Indiana 0.95 40 (1/10) 71 70.8 $\pm$ 0.63 71
socfb-UIllinois 0.94 48 (1/10) 92 92 92
Cit-HepPh 0.98 20 (1/10) 23 23 23
socfb-UF 0.94 69 (1/10) 93 93 93
Email-Enron 0.96 13 (1/10) 26 26 26
socfb-Penn94 0.94 42 (1/10) 63.4 $\pm$ 3.37 63.3 $\pm$ 3.59 65
soc-brightkite 0.97 43 (1/10) 53 53 53
socfb-OR 0.92 24 (1/10) 50.9 $\pm$ 0.32 50.9 $\pm$ 0.32 51
soc-slashdot 0.97 10 (1/10) 35 35 35
wordnet-words 1 29.2 $\pm$ 2.71 (6/10) 32 32 32
witter_combined 0.97 87 (1/10) 99 99 99
G_n_pin_pout 0.5 4 (3/10) 7.9 $\pm$ 0.32 7.9 $\pm$ 0.32 8
yahoo-msg 0.96 24 (1/10) 23.4 $\pm$ 1.84 23.9 $\pm$ 1.1 25
web-sk-2005 0.99 83 (3/10) 84 84 84
soc-douban 0.72 9 (1/10) 23 23 23
wave 0.71 8 (3/10) 14 14 14
web-arabic-2005 1 90.7 $\pm$ 5.36 (10/10) 102 102 102
Loc-Gowalla 1 23.5 $\pm$ 0.71 (2/10) 21 21 29
soc-gowalla 0.99 25 (1/10) 23 23.8 $\pm$ 2.53 31
ca-dblp-2010 1 70.6 $\pm$ 4.86 (10/10) 75 75 75
ca-citeseer 1 81.6 $\pm$ 5.36 (10/10) 87 87 87
coAuthorsCiteseer 1 81.6 $\pm$ 5.36 (10/10) 87 87 87
web-Stanford 0.99 62 (1/10) 64.2 $\pm$ 3.43 59.9 $\pm$ 4.12 67.5 $\pm$ 0.71
coAuthorsDBLP 1 103 $\pm$ 9.7 (10/10) 115 115 115
ca-dblp-2012 1 104 $\pm$ 9.52 (10/10) 114 114 114
com-dblp 1 101.4 $\pm$ 5.76 (10/10) 114 114 114
cnr-2000 1 84 (1/10) 81.9 $\pm$ 1.45 81.8 $\pm$ 1.55 84
web-NotreDame 1 153.5 $\pm$ 1.43 (10/10) 154.3 $\pm$ 0.48 154.2 $\pm$ 0.42 155
ca-MathSciNet 1 23.8 $\pm$ 0.79 (10/10) 24.6 $\pm$ 0.84 24.5 $\pm$ 0.85 25
coPapersDBLP 1 316.9 $\pm$ 23 (10/10) 333 $\pm$ 3.94 333 $\pm$ 3.94 337
auto 0.77 9.5 $\pm$ 0.71 (2/10) 11 11 11
web-it-2004 1 431.7 $\pm$ 0.48 (10/10) 431.8 $\pm$ 0.42 432 432
soc-delicious 0.88 30 (1/10) 31.4 $\pm$ 3.86 31.4 $\pm$ 3.86 37
ca-coauthors-dblp 1 316.9 $\pm$ 23 (10/10) 333.5 $\pm$ 4.12 330.2 $\pm$ 10.01 337
soc-digg 0.9 76 (1/10) 104 $\pm$ 13.7 98.8 $\pm$ 12.56 116.9 $\pm$ 0.32
eu-2005 0.99 387 (1/10) 191.3 $\pm$ 115.84 127.2 $\pm$ 26.92 405.1 $\pm$ 0.32
web-Google 1 40.7 $\pm$ 3.21 (3/10) 33.7 $\pm$ 2.63 37.1 $\pm$ 6.08 44
soc-pokec 0.99 31 (1/10) 24.6 $\pm$ 2.37 23.8 $\pm$ 1.55 25
rgg_n_2_22_s0 0.98 19 (1/10) 20.1 $\pm$ 0.99 20 $\pm$ 0.94 21
rgg_n_2_23_s0 0.99 21 (1/10) 18.3 $\pm$ 0.48 19 $\pm$ 1.41 22
hugetrace-00010 0.5 4 (10/10) N/A N/A 5
hugetrace-00020 0.5 4 (10/10) N/A N/A 5
rgg_n_2_24_s0 0.95 20 (1/10) 20.6 $\pm$ 1.26 20.9 $\pm$ 1.37 25
Effectiveness and Stability.
The results in Table 1 and Table 2 show the superior effectiveness and stability of EDQC. A quantitative summary highlights this advantage. In the first setting (Table 1), across the 47 instances where results differed, EDQC achieves or ties for the best average size in 46 instances. This compares favorably to the other methods, where NuQClqR1 achieved or tied for the best size in 29 instances, followed by NuQClqF1 (28 instances) and NBSim (10 instances). Similarly, in the second setting (Table 2), EDQC achieves or ties for the best size in 68 out of 69 instances. For comparison, the counts for the other algorithms were: NuQClqR1 (46 instances), NuQClqF1 (45 instances), and FastNBSim (2 instances).
Beyond average size, EDQC demonstrates exceptional stability. It consistently reports a standard deviation close to zero (e.g., $0.32$ on the large eu-2005 graph), even on complex instances. This is in stark contrast to the NuQClq variants, which frequently exhibit high variance on the same instances (e.g., a STD of $115.84$ for NuQClqF1 on eu-2005). Furthermore, the results highlight the unreliability of the similarity-based methods. FastNBSim, for instance, often fails to achieve its own best-found density (succeeding only 1/10 times on many graphs). This lack of explicit density control makes them less suitable for applications requiring highly cohesive structures, a task where EDQC’s reliability is a distinct advantage.
Figure 2: Runtime comparison of EDQC against baselines on a log-log scale.
Efficiency.
The efficiency of EDQC is compared against baselines in Figure 2, where, following prior work sunnumds , instances where the runtime of both algorithms is below 0.1 seconds are excluded. On this log-log plot, points falling below the $1\times$, $10\times$, and $100\times$ lines indicate that EDQC is faster than the baseline by the corresponding factor or more. The results clearly show that EDQC is substantially more efficient than the high-performing local search methods, NuQClqF1 and NuQClqR1, with a large number of points falling below the $10\times$ line, signifying a speedup of over an order of magnitude. In comparison to the fast similarity-based heuristics (NBSim and FastNBSim), EDQC’s runtime remains highly competitive, as most of their corresponding points cluster around the $1\times$ line. Overall, this demonstrates that EDQC strikes an excellent balance between achieving state-of-the-art solution quality and maintaining high computational efficiency.
4.5 Statistical Analysis
(a) Comparison in the NBSim setting.
(b) Comparison in the FastNBSim setting.
Figure 3: Critical Difference diagrams for algorithm ranks on all 75 datasets (significance level 0.05). A horizontal bar connects algorithms with no statistically significant difference.
To assess statistical significance, we applied the Friedman test friedman1937use on the algorithm rankings across all 75 datasets. The null hypothesis that all algorithms perform equally was rejected in both settings ($p<0.05$), allowing us to proceed with the post-hoc Nemenyi test demvsar2006statistical to identify pairwise performance differences.
The results are visualized using Critical Difference (CD) diagrams in Figure 3. In these diagrams, algorithms are plotted on an axis according to their average rank across all datasets; a lower rank (further to the left) is better. Groups of algorithms whose performance is not statistically significantly different are connected by a horizontal bar, the length of which corresponds to the critical difference.
The top plot in Figure 3 shows the comparison for the NBSim setting. EDQC achieves the best average rank, and this leading position is statistically significant. The post-hoc test reveals that EDQC’s performance is significantly better than all three other baselines: NuQClqF1, NuQClqR1, and NBSim. A similar trend is observed in the FastNBSim setting, shown in the bottom plot. Once again, EDQC obtains the best average rank. Its superiority is also confirmed to be statistically significant, as it significantly outperforms NuQClqF1, NuQClqR1, and FastNBSim.
In summary, the statistical tests provide conclusive evidence for the superiority of EDQC. In both experimental settings, EDQC not only achieves the best average rank but also demonstrates a statistically significant performance advantage over all compared state-of-the-art baselines. These results provide strong statistical evidence that EDQC represents a reliably superior approach for this task.
4.6 Analysis of EDQC Components
To isolate and verify the contribution of each key component, we conducted an ablation study comparing the full EDQC against five variants across 112 unique experimental settings (consolidating settings where both configurations yielded the same $\gamma$ value). These variants were designed to test the core aspects of our algorithm: the subgraph identification strategy (EDQC-1), the energy-guided ranking (EDQC-2, EDQC-3), the refinement phase (EDQC-4), and the nature of the diffusion process itself (EDQC-5).
The results, summarized in Table 3, provide conclusive evidence that each component is essential for EDQC’s superior performance. The most dramatic performance degradation occurred with EDQC-5, which replaced our adaptive diffusion with a naive random process, and EDQC-1, which substituted conductance-based identification with a max-density heuristic. The full EDQC algorithm proved superior in 106 and 69 instances, respectively. These results underscore that our principled diffusion and extraction mechanisms form the foundational pillars of the algorithm’s success. The importance of the energy-guided ranking was also clearly demonstrated; disrupting the descending energy order with a random (EDQC-2) or ascending (EDQC-3) ranking severely hampered performance, leading to 56 and 70 wins for the full EDQC. Finally, disabling the refinement phase (EDQC-4) also resulted in a noticeable performance drop in 16 instances, confirming that this final step is vital for optimizing the solution size.
Additionally, we confirmed the robustness of our synchronous design. As variants with a fixed processing order for active vertices produced identical results to the full EDQC, the order-independence guaranteed by the synchronous design was empirically verified.
Table 3: Comparison of the full EDQC with its five simplified variants across 112 experimental settings. #bet and #wor represent the number of instances where EDQC achieves better and worse quasi-clique sizes, respectively.
Variant Description #bet #wor
(EDQC ¿ Variant) (EDQC ¡ Variant)
EDQC-1 No Conductance 69 0
EDQC-2 Random Ranking 56 1
EDQC-3 Ascending Ranking 70 1
EDQC-4 No Refinement Phase 16 1
EDQC-5 Naive Random Diffusion 106 0
Figure 4: Subgraph density vs. total retained energy on eight representative datasets. Each blue dot represents one of 1,000 randomly sampled subgraphs of a fixed size; the red marks the quasi-clique of the same size found by EDQC.
4.7 Empirical Validation of the Energy-Density Premise
To empirically validate our algorithm’s core premise that energy concentrates in dense regions, we analyzed the relationship between subgraph density and energy retention on eight representative datasets. For each dataset, we compared the quasi-clique found by EDQC against 1,000 randomly sampled subgraphs of the same size. The results, visualized in Figure 4, reveal a strong positive correlation between density and retained energy, confirmed by high Pearson correlation coefficients (typically >0.90>0.90). More importantly, the quasi-clique discovered by EDQC (red marks) consistently appears as a top-right outlier, indicating that it captures a disproportionately high amount of energy even among other subgraphs of similar density. This strong alignment provides the empirical justification for using the energy ranking as a powerful guide in the discovery of quasi-cliques.
4.8 Parameter Sensitivity
This section provides the detailed results of our parameter sensitivity analysis, which was designed to evaluate the impact of the number of diffusion rounds $T$ and the activation threshold $\theta$ on EDQC’s performance. For this study, we selected three representative datasets to cover different scales: ego-facebook (a small graph, $|V|\leq 10^{4}$), rgg_n_2_19_s0 (a medium graph, $10^{4}<|V|\leq 10^{6}$), and web-Google (a large graph, $|V|>10^{6}$).
On each of these graphs, we evaluated EDQC by testing 15 combinations of $T\in\{1,2,3\}$ and $\theta\in\{10^{-4},5\times 10^{-4},10^{-3},5\times 10^{-3},10^{-2}\}$. This evaluation was repeated under three different density thresholds, $\gamma\in\{0.7,0.8,0.9\}$, to ensure a comprehensive analysis. We report both the discovered quasi-clique size and the runtime for each configuration. The results for the small, medium, and large graphs are presented in Figure 57.
Figure 5: Parameter sensitivity analysis for EDQC on the small representative graph (ego-facebook). Each of the three rows corresponds to a different density threshold $\gamma$. For each $\gamma$, the left plot shows the resulting quasi-clique size versus the activation threshold $\theta$ (on a log scale), and the right plot shows the corresponding runtime.
Figure 6: Parameter sensitivity analysis on the medium representative graph (rgg_n_2_19_s0). The layout of the plots is identical to that described in the caption of Figure 5.
Figure 7: Parameter sensitivity analysis on the large representative graph (web-Google). The layout of the plots is identical to that described in the caption of Figure 5.
Across all three datasets, a consistent pattern emerges. In terms of solution quality (size), configurations with $T=2$ and $T=3$ consistently and significantly outperform $T=1$, especially for smaller values of the activation threshold ($\theta\leq 10^{-3}$). As $\theta$ increases, making vertex activation stricter, the performance of all settings tends to decline. Regarding efficiency, the runtime plots clearly show that the computational cost increases with $T$. While $T=3$ occasionally finds a marginally larger quasi-clique than $T=2$, its runtime is substantially higher across all settings. For instance, in the large graph setting (Figure 7), the runtime for $T=3$ can be several times that of $T=2$ for little to no improvement in solution size.
Based on this analysis, we conclude that the configuration $T=2$ provides the best trade-off between solution quality and computational cost. Paired with a small activation threshold such as $\theta=10^{-4}$, which consistently lies in the stable, high-performance region shown in the plots, this configuration represents an effective and efficient choice. This justifies our use of $T=2$ and $\theta=10^{-4}$ for all main experiments presented in the paper.
5 Conclusion
In this paper, we introduced EDQC, a novel algorithm for quasi-clique discovery designed to mitigate key limitations of existing methods, such as their sensitivity to random seeds and lack of explicit edge-density guarantees. We demonstrated that the effectiveness of EDQC stems from its two-stage design. First, an adaptive energy diffusion process generates an energy ranking that highlights structurally cohesive regions. Then, guided by this ranking, a principled extraction mechanism based on conductance minimization identifies and refines a high-quality subgraph. To the best of our knowledge, EDQC is the first method to incorporate energy diffusion into this task, offering a robust alternative to traditional heuristics.
Our extensive experiments on 75 real-world graphs confirmed the effectiveness of this approach. The results demonstrated that EDQC finds larger quasi-cliques on most datasets, with consistently lower variance across runs, highlighting its superior stability. Furthermore, statistical tests confirmed the significance of these findings, and our analysis showed that EDQC maintains a competitive runtime, striking an excellent balance between solution quality and efficiency.
Future work can extend the EDQC framework in several promising directions. Generalizing the diffusion model for weighted and attributed graphs is an important next step. Applying EDQC to downstream tasks, such as anomaly detection in financial networks and community detection in social graphs, could further validate its practical utility. Finally, exploring more advanced refinement strategies may lead to further improvements in solution quality.